# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v3

@@ A.1 Unforgeability Proof (Full)
    Simulate honest parties’ messages $\mathbf{W}_{j},\mathbf{U}_{j}$ as uniform (justified by Lemma 4.6)
Analysis: By Lemma 4.6, each honest party’s masked value is computationally indistinguishable from uniform. Hybrid over $|S\setminus C|$ honest parties: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq|S|\cdot\epsilon_{\mathsf{PRF}}$.
@@
with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),\tau)<2\gamma_{1}$, solving M-SIS.
@@ A.2 Privacy Proof (Full)
Round 1 Simulation:
  1. 1.
    For each honest $j\in S\setminus C$: sample $\mathsf{Com}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ uniformly
  2. 2.
    Output $\{\mathsf{Com}_{j}\}_{j\in S\setminus C}$
Round 2 Simulation:
  1. 1.
    Expand challenge $c$ from $\tilde{c}$
  2. 2.
    Compute $\mathbf{w}_{1}=\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})$
  3. 3.
    Recover $\mathbf{w}$ as any vector with $\mathsf{HighBits}(\mathbf{w},2\gamma_{2})=\mathbf{w}_{1}$
  4. 4.
    Sample $\mathbf{W}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$ for $j\in S\setminus C$ subject to $\sum_{j\in S\setminus C}\mathbf{W}_{j}=\mathbf{w}-\sum_{i\in C}\mathbf{W}_{i}$
  5. 5.
    Output $\{(\mathbf{W}_{j},r_{j})\}_{j\in S\setminus C}$ with random $r_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
Round 3 Simulation:
  1. 1.
    Receive corrupted responses $\{\mathbf{U}_{i}\}_{i\in C}$ from $\mathcal{A}$
  2. 2.
    Sample $\mathbf{U}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ for $j\in S\setminus C\setminus\{j^{*}\}$ uniformly
  3. 3.
    Set $\mathbf{U}_{j^{*}}=\mathbf{z}-\sum_{i\in C}\mathbf{U}_{i}-\sum_{j\neq j^{*}}\mathbf{U}_{j}$ for arbitrary honest $j^{*}$
  4. 4.
    Output $\{\mathbf{U}_{j}\}_{j\in S\setminus C}$
Hybrid 0 (Real).
Real protocol execution.
Hybrid 1.
Replace PRF outputs for honest-honest seed pairs with truly random values. For each pair $(j,k)$ with $j,k\in S\setminus C$ and $j<k$: replace $\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})$ with $\mathbf{r}_{j,k}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{d}$. Analysis: By PRF security and hybrid argument over $\binom{|S\setminus C|}{2}$ pairs: $|\Pr[\mathsf{H}_{1}]-\Pr[\mathsf{H}_{0}]|\leq\binom{|S\setminus C|}{2}\cdot\epsilon_{\mathsf{PRF}}$.
Hybrid 2 (Simulated).
Replace honest masked values with simulator’s output. Analysis: In $\mathsf{H}_{1}$, honest party $j$’s mask includes $\mathbf{r}_{j,k}$ for some honest $k\neq j$ (exists since $|S\setminus C|\geq 2$). This term is uniform in $R_{q}^{d}$ and unknown to $\mathcal{A}$. Thus:
$\mathbf{U}_{j}=(\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\text{known terms})+\mathbf{r}_{j,k}\equiv\mathsf{Uniform}(R_{q}^{\ell})$
The simulator’s uniform sampling produces identical distribution: $\Pr[\mathsf{H}_{2}]=\Pr[\mathsf{H}_{1}]$.
Appendix B Distributed Key Generation Protocol
B.1 Protocol Description
The DKG proceeds in two phases:
@@ B.1 Protocol Description
The final secret is $\mathbf{s}_{1}=\sum_{i=1}^{N}\mathbf{s}_{1}^{(i)}$, which no individual party knows.
B.2 Security Properties
  • Correctness: Honest execution produces a valid key pair.
  • Secrecy: Any coalition of <T<T parties learns nothing about $\mathbf{s}_{1}$ beyond what is revealed by the public key (follows from Shamir’s threshold security).
@@ Appendix D Numerical Verification of Rényi Bounds
For signing set size $|S|=n$, each party samples nonces from range $N=2\gamma_{1}/n$. Using Lemma 4.2 of [Raccoon2024] with shift $c=\beta$:
@@
Table 12: Numerical Rényi divergence bounds for ML-DSA-65. All values computed using proven Lemma 4.2 of [Raccoon2024].
The multiplicative security loss $R_{2}^{\epsilon}$ remains below $1.05$ even for $|S|=33$ signers.
@@ Appendix G Nonce Range Rounding Details
Each party samples $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$. When $\gamma_{1}$ is not divisible by $|S|$, the sum $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ has range $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor,|S|\cdot\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$, slightly smaller than standard $\gamma_{1},\gamma_{1}]^{n\ell}$.
@@ H.2 Hybrid Functionalities
Statistical distance: The bias is $(2^{23}-q)/2^{23}=8191/8388608<2^{-10}$. For $m=1536$ coefficients, the total statistical distance is at most $m\cdot 2^{-10}<2^{1}$ by union bound, which is not negligible.
@@ Appendix I Complete Ideal Functionality for Profile P2
Theorem M.1 (UC Security of Profile P2).
@@ 5 Output
    If $|\mathsf{responses}|\geq T+1$: trigger 2PC phase
@@
    Select $T+1$ signers with available pre-signing sets
@@
    Aggregate $\mathbf{w}=\sum_{i}\mathbf{w}_{i}$
@@
    Retrieve $(\mathbf{y}_{i},\mathbf{w}_{i})$ for assigned $\mathsf{set\_id}$
@@
    Compute Lagrange coefficient $\lambda_{i}$
@@
    Compute:
    $\mathbf{U}_{i}$ $=\mathbf{y}_{i}+\lambda_{i}c\mathbf{s}_{1,i}+\mathbf{m}_{i}^{(s1)}$
    $\mathbf{V}_{i}$ $=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$
@@
    Wait for $T+1$ responses (or timeout)
@@
    Aggregate: $\mathbf{U}=\sum_{i}\mathbf{U}_{i}$ (masks cancel)
@@
    If $\mathsf{result}=1$: CPs compute hints, combiner outputs signature
@@
      Sample random $\tilde{\mathbf{U}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$
@@
      Send $(\tilde{\mathbf{U}}_{i},\tilde{\mathbf{V}}_{i})$ to adversary (via combiner)
@@
    Justification: Masks $\mathbf{m}_{i}^{(s1)},\mathbf{m}_{i}^{(s2)}$ are uniform one-time pads. Real $\mathbf{U}_{i},\mathbf{V}_{i}$ are uniform; simulated values have same distribution.
@@
$\mathbf{U}_{i}$ $=\mathbf{y}_{i}+\lambda_{i}c\mathbf{s}_{1,i}+\mathbf{m}_{i}^{(s1)}$
@@
The masks $\mathbf{m}_{i}^{(s1)},\mathbf{m}_{i}^{(s2)}$ are derived from pairwise PRF seeds shared only with other honest signers. Against an adversary controlling fewer than all honest signers:
  • At least one honest signer contributes an unknown mask component
  • This mask is uniform and independent of the rest
  • By one-time pad: $\mathbf{U}_{i},\mathbf{V}_{i}$ are uniform
Simulated $\tilde{\mathbf{U}}_{i},\tilde{\mathbf{V}}_{i}$ are sampled uniformly, matching the real distribution.
@@
    Privacy: Honest signer shares hidden if $\geq 2$ honest signers
@@
    Liveness: Completes if $T+1$ signers respond within $\Delta$ and both CPs available
@@
EUF-CMA: Pre-signing doesn’t change the signature distribution. The commitment binding ensures signers commit to nonces before seeing challenges. Reduction to Module-SIS identical to P3.
@@
Privacy: By Lemma P.9, masked responses are uniformly distributed. The 2PC reveals only the result bit. Combining: adversary view simulatable from public values + corrupted shares + result bit.
@@
    Expected retries: 3.4 ($\approx$29% success rate)
@@
    Expected total rounds: 6.8 (2 rounds $\times$ 3.4 attempts)
@@
Total: $2\times 3.4\approx 6.8$ expected rounds (vs. $5\times 3.4=17$ for P2).
@@
P3 (original) 5 3.5 17.5
P3+ 2 3.5 7
Table 16: Round complexity comparison across profiles.
@@
Lemma Q.3 (Mask Hiding – Formal Statement).
Let $C\subsetneq S$ be the set of corrupted parties in signing set $S$, with $|S\setminus C|\geq 2$. For any honest party $j\in S\setminus C$, the masked value:
$\mathbf{U}_{j}=\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}$
is computationally indistinguishable from uniform in $R_{q}^{\ell}$, given the adversary’s view consisting of:
  • Corrupted shares $\{\mathbf{s}_{1,i},\mathbf{s}_{2,i}\}_{i\in C}$
  • Corrupted seeds $\{\mathsf{seed}_{i,k}\}_{i\in C,k\in S}$
  • All protocol messages
@@
Decompose $\mathbf{m}_{j}$ based on counterparty corruption status:
$\mathbf{m}_{j}$ $=\underbrace{\sum_{k\in C\cap S:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in C\cap S:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\mathbf{m}_{j}^{\text{known}}:\text{ computable by }\mathcal{A}}$
@@
$\quad+\underbrace{\sum_{k\in(S\setminus C)\setminus\{j\}:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in(S\setminus C)\setminus\{j\}:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\mathbf{m}_{j}^{\text{unknown}}:\text{ unknown to }\mathcal{A}}$
Since $|S\setminus C|\geq 2$, there exists at least one honest party $k^{\prime}\in(S\setminus C)\setminus\{j\}$.
Claim: $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ (assuming WLOG $j<k^{\prime}$) is computationally indistinguishable from uniform in $R_{q}^{\ell}$.
Proof of Claim:
  1. 1.
    The seed $\mathsf{seed}_{j,k^{\prime}}$ is shared only between honest parties $j$ and $k^{\prime}$
  2. 2.
    $\mathcal{A}$ does not know $\mathsf{seed}_{j,k^{\prime}}$ (parties $j,k^{\prime}\notin C$)
  3. 3.
    The domain separator $\mathsf{dom}$ includes a fresh nonce, so this is a fresh PRF evaluation
  4. 4.
    By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ is indistinguishable from $\mathbf{r}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$
Conclusion:
$\mathbf{U}_{j}$ $=\underbrace{\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}^{\text{known}}}_{\text{fixed from }\mathcal{A}\text{'s view}}+\underbrace{\mathbf{m}_{j}^{\text{unknown}}}_{\text{contains uniform-looking term}}$
Since $\mathbf{m}_{j}^{\text{unknown}}$ contains at least one term indistinguishable from uniform, and uniform plus fixed equals uniform:
$\mathbf{U}_{j}\stackrel{{\scriptstyle c}}{{\approx}}\mathsf{Uniform}(R_{q}^{\ell})$
The distinguishing advantage is bounded by the PRF advantage:
$\mathsf{Adv}^{\text{mask-hiding}}\leq|S\setminus C|\cdot\epsilon_{\mathsf{PRF}}$
where we use a hybrid over all honest-honest PRF terms.

### v4

@@ A.1 Unforgeability Proof (Full)
    Simulate honest parties’ responses $\mathbf{z}_{j}$ by sampling from a high-entropy distribution (justified by Theorem 4.4: conditional min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits) and their masked values $\mathbf{W}_{j},\mathbf{V}_{j}$ as uniform (justified by Lemma 4.9)
Analysis: By Theorem 4.4, each honest party’s response $\mathbf{z}_{j}$ has high conditional min-entropy (statistical, no computational assumptions). The masked commitment and r0-check values $\mathbf{W}_{j},\mathbf{V}_{j}$ are computationally indistinguishable from uniform by Lemma 4.9. Hybrid over $|S\setminus C|$ honest parties: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq n\ell\cdot(1-2\gamma_{1}/(|S|\cdot q))+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (the first term from response simulation is negligible for practical $|S|$; the second from mask-based values uses the union bound over all $\binom{N}{2}$ pairwise PRF keys).
@@
with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),2)<2\gamma_{1}$, solving M-SIS.
@@ A.2 Privacy Proof (Full)
With the Shamir nonce DKG, privacy of the response $\mathbf{z}_{h}$ is statistical (no computational assumptions) and does not require a hybrid argument. The proof simplifies considerably compared to the computational (PRF-based) approach.
@@
Round 1 Simulation: For each honest $j\in S\setminus C$, sample $\mathsf{Com}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ uniformly.
[…]
Round 3 Simulation: For each honest $j\in S\setminus C$, sample $\mathbf{z}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ uniformly. This is justified by Theorem 4.4: the real $\mathbf{z}_{h}$ has high conditional min-entropy ($\geq 5\times$ the secret key entropy for $|S|\leq 17$) from the adversary’s perspective, so the simulator’s uniform sampling introduces negligible statistical distance. The $\mathbf{V}_{j}$ values are simulated uniformly by the same PRF-based argument as for $\mathbf{W}_{j}$.
Indistinguishability.
The simulation is statistical for $\mathbf{z}_{j}$ (high min-entropy, negligible statistical distance from uniform) and computationally indistinguishable for $\mathbf{W}_{j}$ (where $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$) and $\mathbf{V}_{j}$ (PRF security). The total distinguishing advantage is $n\ell\cdot(1-2\gamma_{1}/(|S|\cdot q))+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$, where the first term (negligible for practical $|S|$) arises from the response simulation and the second from mask-based values. In particular, no PRF assumption is needed for response privacy.
Appendix B Distributed Key Generation Protocols
Our protocol uses two structurally similar but functionally distinct DKG protocols: a key DKG (executed once during setup) and a nonce DKG (executed before each signing session, typically preprocessed offline). Both follow the Feldman VSS paradigm but differ in their input distributions and security requirements.
B.1 Key DKG (One-Time Setup)
The key DKG produces a degree-(T1)(T\!-\!1) Shamir sharing of the long-term signing key $(\mathbf{s}_{1},\mathbf{s}_{2})$, replacing the trusted dealer assumption.
@@ B.1 Key DKG (One-Time Setup)
The final secret is $\mathbf{s}_{1}=\sum_{i=1}^{N}\mathbf{s}_{1}^{(i)}$, which no individual party knows. For lattice-specific VSS/DKG, see [FS24].
B.2 Nonce DKG (Per-Session Preprocessing)
The nonce DKG produces a fresh degree-(T1)(T\!-\!1) Shamir sharing of the signing nonce $\mathbf{y}$ for each signing session. Unlike the key DKG, the nonce DKG samples higher-degree coefficients uniformly from $R_{q}$ (rather than from a short-coefficient distribution), which is critical for achieving statistical share privacy.
The protocol is specified formally in Algorithm 1 of the main paper (Section 3.4). Each party $i\in S$ generates a degree-(T1)(T\!-\!1) polynomial $\hat{f}_{i}(X)$ whose constant term $\hat{\mathbf{y}}_{i}^{(0)}$ is sampled from the reduced nonce range $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ and whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$. The resulting aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}^{(0)}$ follows the Irwin-Hall distribution analyzed in Theorem 4.11, while each honest party’s individual share $\mathbf{y}_{h}=\sum_{i\in S}\hat{f}_{i}(h)$ has high conditional min-entropy from the adversary’s perspective (Theorem 4.4).
B.3 Security Properties
Both protocols share the following properties:
  • Correctness: Honest execution produces a valid sharing of the intended secret.
  • Secrecy: Any coalition of <T<T parties learns nothing about the secret beyond what is revealed by public commitments (follows from Shamir’s threshold security).
@@ B.3 Security Properties
The nonce DKG additionally provides statistical privacy (no computational assumptions) of individual shares (Theorem 4.4), which is the key enabler for eliminating the two-honest requirement for response privacy in coordinator-based profiles (P1, P3+; see Remark 3.7). The key DKG does not require this property, since key shares are never revealed during signing.
@@ Appendix D Numerical Verification of Rényi Bounds
For signing set size $|S|=n$, each party samples nonces from range $N=2\lfloor\gamma_{1}/n\rfloor+1$. For readability, we use the approximation $N\approx 2\gamma_{1}/n$ (the difference is at most $1/N<2\times 10^{-5}$ for all relevant $n$, and the exact discrete range is used in the implementation). Using Lemma 4.2 of [Raccoon2024] with $\alpha=2$ (where $\alpha(\alpha-1)/2=1$) and shift $c=\beta$:
@@
Table 12: Per-coordinate Rényi divergence bounds for ML-DSA-65. All values computed using proven Lemma 4.2 of [Raccoon2024]. Full-vector: $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$.
The per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}$ remains below $1.05$ for $|S|\leq 33$. The full-vector divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ ranges from $\approx 2^{5.4}$ ($|S|=17$) to $\approx 2^{75}$ ($|S|=33$); see Theorem 4.11 for the impact on the security bound.
@@ Appendix G Nonce Range Rounding Details
Each party’s constant-term contribution in the nonce DKG is $\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$. When $\gamma_{1}$ is not divisible by $|S|$, the nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ has range $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor,|S|\cdot\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$, slightly smaller than standard $\gamma_{1},\gamma_{1}]^{n\ell}$.
@@ H.2 Hybrid Functionalities
Per-coefficient bias: $(2^{23}-q)/2^{23}=8191/8388608<2^{-10}$. A naïve union bound over $m=1536$ coefficients gives $m\cdot 2^{-10}\approx 1.5$, which exceeds 1 and is therefore vacuous (statistical distance is at most 1 by definition).
@@ Appendix I Complete Ideal Functionality for Profile P2
Theorem M.1 (UC Security of Profile P2 (Full Proof of Theorem 6.2)).
@@ 5 Output
    If $|\mathsf{responses}|\geq T$: trigger 2PC phase
@@
    Select $T$ signers with available pre-signing sets
@@
    Aggregate $\mathbf{w}=\sum_{i}\lambda_{i}\cdot\mathbf{w}_{i}$ $\triangleright$ Lagrange reconstruction
@@
    Retrieve nonce DKG share $\mathbf{y}_{i}$ and commitment $\mathbf{w}_{i}$ for assigned $\mathsf{set\_id}$
@@
    Compute:
    $\mathbf{z}_{i}$ $=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$
    $\mathbf{V}_{i}$ $=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$
    where $\lambda_{i}$ is the Lagrange coefficient for party $i$ in signing set $S$
@@
    Send $(\mathbf{z}_{i},\mathbf{V}_{i})$ to combiner
@@
    Wait for $T$ responses (or timeout)
@@
    Aggregate response: $\mathbf{z}=\sum_{i\in S}\lambda_{i}\cdot\mathbf{z}_{i}$ (Lagrange reconstruction)
@@
    If $\mathsf{result}=1$: The 2PC circuit also outputs the hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$ (computed inside the circuit using $c\mathbf{s}_{2}=\mathbf{S}_{1}+\mathbf{S}_{2}$); combiner outputs signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
@@
      Sample random $\tilde{\mathbf{z}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$
@@
      Send $(\tilde{\mathbf{z}}_{i},\tilde{\mathbf{V}}_{i})$ to adversary (via combiner)
@@
    Justification: By Theorem 4.4 (statistical privacy), the Shamir nonce DKG ensures each honest party’s response $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ has high conditional min-entropy ($\geq 5\times$ secret key entropy for $|S|\leq 17$) from the adversary’s perspective without computational assumptions. The $\mathbf{V}_{i}$ values are uniform by PRF-based mask hiding (Lemma 4.9). Simulated (uniform) values have negligible statistical distance from real values.
@@
$|\Pr[\mathsf{H}_{3}=1]-\Pr[\mathsf{H}_{4}=1]|\leq\mathsf{negl}(\kappa)$ (statistical indistinguishability)
@@
$\mathbf{z}_{i}$ $=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$
@@
By Theorem 4.4 (statistical privacy), the Shamir nonce DKG ensures that the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) nonce polynomial, leaving one degree of freedom that provides high conditional min-entropy. Due to the bounded-nonce constraint, $\mathbf{z}_{i}$ is not perfectly $R_{q}^{\ell}$-uniform but has min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits, vastly exceeding the secret key entropy. The $\mathbf{V}_{i}$ values contain pairwise masks $\mathbf{m}_{i}^{(s2)}$ derived from PRF seeds shared only with other honest signers; against an adversary controlling fewer than all honest signers, at least one honest signer contributes an unknown mask component that acts as a one-time pad.
Simulated $\tilde{\mathbf{z}}_{i},\tilde{\mathbf{V}}_{i}$ are sampled uniformly, matching the real distribution.
@@
    Privacy: Response shares $\mathbf{z}_{i}$ enjoy statistical privacy (Theorem 4.4) even with a single honest signer; commitment and r0-check values hidden by pairwise masks if $\geq 2$ honest signers
@@
    Liveness: Completes if $T$ signers respond within $\Delta$ and both CPs available
@@
EUF-CMA: Pre-signing does not change the signature distribution. The commitment binding ensures signers commit to nonces before seeing challenges. Reduction to Module-SIS identical to P3.
@@
Privacy: By Lemma P.9, honest signers’ responses have high conditional min-entropy (statistical, via Shamir nonce DKG). The $\mathbf{V}_{j}$ values are computationally hidden by pairwise masks. The 2PC reveals only the result bit. Combining: adversary view simulatable from public values + corrupted shares + result bit.
@@
    Expected retries: 3.7 ($\approx$27% success rate)
@@
    Expected total rounds: 7.4 (2 rounds $\times$ 3.7 attempts)
@@
Total: $2\times 3.7\approx 7.4$ expected rounds (vs. $5\times 3.5=17.5$ for P2).
@@
P3 5 3.5 17.5
P3+ 2 3.7 7.4
Table 16: Round complexity comparison across profiles. P3 denotes the synchronous 2PC variant (5 rounds) prior to the semi-async optimization; P3+ is the final semi-async protocol presented in Section 6.5.
@@
Theorem Q.3 (Statistical Share Privacy – Formal Statement).
Let $S$ be a signing set with $|S|\geq T$, let $C\subset S$ with $|C|\leq T-1$ be the set of corrupted parties, and let $h\in S\setminus C$ be an honest party. In the Shamir nonce DKG protocol, the adversary’s view of $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has conditional min-entropy:
$H_{\infty}\!\left(\mathbf{z}_{h}\mid\{\mathbf{f}_{j}\}_{j\in C},\;\{\mathbf{f}_{h}(j)\}_{j\in C},\;c\right)\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
No computational assumptions are needed. Note that the conditioning does not include individual $\mathbf{W}_{i}=\lambda_{i}\mathbf{A}\mathbf{y}_{i}+\mathbf{m}_{i}^{(w)}$ values, which are sent only to the coordinator (not broadcast) in the protocol. Including the coordinator’s broadcast $\mathbf{w}_{1}=\mathsf{HighBits}(\sum_{i}\mathbf{W}_{i},2\gamma_{2})$ does not reduce the bound, since $\mathbf{w}_{1}$ is already implicit in any valid signature. For ML-DSA-65 with $|S|\leq 17$, this exceeds 24,00024{,}000 bits—over $5\times$ the secret key entropy $H(\mathbf{s}_{1,h})\leq 4058$ bits.
@@
The proof proceeds in four steps, following a null space argument on the honest party’s polynomial.
The adversary’s knowledge about the honest party’s polynomial $\mathbf{f}_{h}$ consists of the evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ received during the DKG. Since $|C|\leq T-1$ and $\mathbf{f}_{h}$ has degree T1T-1 (hence $T$ coefficients), the adversary has at most T1T-1 linear equations in $T$ unknowns.
Step 1: Null space. The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space containing the polynomial $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$, which has degree $|C|\leq T-1$ and leading coefficient 1. The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant through the known points and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
Step 2: Posterior constraint on $\mathbf{t}$. The coefficient $\mathbf{a}_{h,T-1}$ is sampled uniformly from $R_{q}^{\ell}$ a priori. However, the constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$. Since $\mathbf{n}(0)\neq 0\pmod{q}$, this constrains each coefficient of $\mathbf{t}$ to at most $2\lfloor\gamma_{1}/|S|\rfloor+1$ values. The posterior distribution of $\mathbf{t}$ is not $\mathbb{Z}_{q}$-uniform, but retains per-coefficient support of this size.
Step 3: Min-entropy at honest point. The honest party’s share is $\mathbf{f}_{h}(x_{h})=\mathbf{g}(x_{h})+\mathbf{t}\cdot\mathbf{n}(x_{h})$. Since $x_{h}\notin C$ and $q$ is prime, $\mathbf{n}(x_{h})\neq 0\pmod{q}$, so multiplication by $\mathbf{n}(x_{h})$ is a bijection on $\mathbb{Z}_{q}$. The evaluation $\mathbf{f}_{h}(x_{h})$ inherits the same per-coefficient support size $2\lfloor\gamma_{1}/|S|\rfloor+1$. The nonce share $\mathbf{y}_{h}=\sum_{j\in C}\mathbf{f}_{j}(x_{h})+\mathbf{f}_{h}(x_{h})$ shifts by a known constant, preserving support.
Step 4: Response. The response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ shifts again by a fixed quantity, yielding $H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\geq n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$.
Lemma Q.5 (Mask Hiding – Formal Statement (for $\mathbf{V}_{i}$ and $\mathbf{W}_{i}$)).
Let $C\subsetneq S$ be the set of corrupted parties in signing set $S$, with $|S\setminus C|\geq 2$. For any honest party $j\in S\setminus C$, the masked r0-check share $\mathbf{V}_{j}=\lambda_{j}c\mathbf{s}_{2,j}+\mathbf{m}_{j}^{(s2)}$ and masked commitment $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$ are computationally indistinguishable from uniform, given the adversary’s view.
Note: This lemma applies to the commitment and r0-check values, which still use pairwise-canceling masks. The response $\mathbf{z}_{j}=\mathbf{y}_{j}+c\cdot\mathbf{s}_{1,j}$ achieves statistical privacy via the Shamir nonce DKG (Theorem Q.3) and does not require masks.
Proof Q.6.
@@
We prove the result for $\mathbf{V}_{j}$; the argument for $\mathbf{W}_{j}$ is identical. Decompose the mask $\mathbf{m}_{j}^{(s2)}$ based on counterparty corruption status:
$\mathbf{m}_{j}^{(s2)}$ $=\underbrace{\sum_{k\in C\cap S:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in C\cap S:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\text{computable by }\mathcal{A}}$
@@
$\quad+\underbrace{\sum_{k\in(S\setminus C)\setminus\{j\}:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in(S\setminus C)\setminus\{j\}:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\text{unknown to }\mathcal{A}}$
Since $|S\setminus C|\geq 2$, there exists at least one honest party $k^{\prime}\in(S\setminus C)\setminus\{j\}$ whose PRF seed $\mathsf{seed}_{j,k^{\prime}}$ is unknown to $\mathcal{A}$. By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ is computationally indistinguishable from uniform in $R_{q}^{k}$. Since the mask contains at least one such uniform-looking term, and uniform plus fixed equals uniform, $\mathbf{V}_{j}\stackrel{{\scriptstyle c}}{{\approx}}\mathsf{Uniform}(R_{q}^{k})$. The distinguishing advantage is bounded by $|S\setminus C|\cdot\epsilon_{\mathsf{PRF}}$.